OpenStudy (anonymous):

This also i can't solve... Two vehichles p and Q,moving at the same speed are acted upon by equal braking forces until they stop.If the mass of P is four times that of Q,the ratio of the distance travelled by P to that travelled by Q is? @Abhisar

OpenStudy (abhisar):

Do u have the answer key ?

OpenStudy (anonymous):

answer key?

OpenStudy (anonymous):

don't know...

OpenStudy (abhisar):

Ok let's see, \(\sf F=ma\), now if force is same or constant then \(\sf a\propto \frac{1}{m}\)

OpenStudy (abhisar):

\(\sf v^2=u^2-2as\), since the vehicles are braked until they stop, there final velocity v will be 0. Now, \(\large \sf s=\frac{u^2}{2a}\)

OpenStudy (abhisar):

So, \(\huge \sf s_p=\frac{u^2}{8a}, s_q=\frac{u^2}{2a}\)

OpenStudy (abhisar):

Where, S\(_p~and~S_q\) are the distance travelled by p and q. Now can u find the ratio ?

OpenStudy (anonymous):

yes...4:1

OpenStudy (abhisar):

\(\huge\checkmark\)

OpenStudy (anonymous):

Thank you...

OpenStudy (vincent-lyon.fr):

Solving with energy is faster. Since initial KE ratio is 4:1, then work done by the friction force must also be 4:1. Since the magnitude of the force is the same, then the distance it must act is also in a 4:1 ratio.

OpenStudy (anonymous):

ok...got it...

OpenStudy (abhisar):

\(\huge\sf\color{green}{\text{✌゚\(\ddot\smile\) ✌゚}}\)